On a buying trip in Los Angeles, Rosaria Perez ordered 120 pieces of jewelry: a number of bracelets at $8 each and a number of necklaces at $11 each. She wrote a check for $1140 to pay for the order. How many bracelets and how many necklaces did Rosaria purchase?
Answer by ikleyn(52754)

.
On a buying trip in Los Angeles, Rosaria Perez ordered 120 pieces of jewelry: a number of bracelets at $8 each and a number
of necklaces at $11 each. She wrote a check for $1140 to pay for the order. How many bracelets and how many necklaces did Rosaria purchase?
~~~~~~~~~~~~~~~~~~~~~~~

 B +   N =  120,       (1)
8B + 11N = 1140.       (2)

From (1), express B = 120 - N and substitute it into (2) replacing B. You will have

8*(120-N) + 11N = 1140,

960 - 8N + 11N = 1140,

3N = 1140 - 960  --->  3N = 180   --->  N = 60.

Then B = 120 - 60 = 60.

Answer.  60 bracelets and 60 necklaces.